OpenStudy (anonymous):

the plane that contains the line x=3t , y=1+t , z=2t and is parallel to the intersection of the planes y+z=-1 and 2x -y+z=0.

OpenStudy (anonymous):

first find the point of intersection of planes then assume plane ax+by+cz+d=0 this point will satisfy the equation. and the three lines willl be perpendicular to normal of the plane so u will get 4 equation and four variables now you can solve

OpenStudy (anonymous):

in detail please

ganeshie8 (ganeshie8):

say, \(a(x-x_1) + b(y-y_1 ) + c(z-z_1) = 0 \) is the equation of required plane

ganeshie8 (ganeshie8):

since it contains the line, x=3t , y=1+t , z=2t and since normal is perpendicular to every line on the plane, we have : \(3a + b + 2c = 0\)

ganeshie8 (ganeshie8):

Also, since the plane is parallel to line of intersection of plane y+z=-1 and 2x -y+z=0 : direction vector of this line of intersection = \(\left<0, 1, 1\right> \times \left<2, -1, 1\right> = \left<1, 1, -1\right>\) the normal of required plane is perpendicular to this direction vector also : \(\large a +b -c = 0\)

ganeshie8 (ganeshie8):

solving above two equations for a, b, c gives : (a, b, c) = (-3, 5, 2)

ganeshie8 (ganeshie8):

so the equation of required plane is : \(-3(x-x_1) + 5(y-y_1 ) + 2(z-z_1) = 0\)

ganeshie8 (ganeshie8):

plugin the point (0, 1, 0) : \(-3(x-0) + 5(y-1 ) + 2(z-0) = 0\)

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

let me knw if smthng doesnt make sense...

OpenStudy (anonymous):

okay thanks again :)

ganeshie8 (ganeshie8):

np :)